OpenStudy (anonymous):

I need help with a calculus problem. The equation for the general form of the circle is 2x^2 +8x +2y^2 -y -2 = 0 . I have found the center (h,k) to equal (-2,1/4) but I can not figure out the radius. Can someone please tell me what the radius is and how to get to it

OpenStudy (anonymous):

@ganeshie8 is there anyway you could help me?

OpenStudy (fibonaccichick666):

So first, can you tell me the standard eq for a circle?

OpenStudy (anonymous):

\[2x ^{2} +8x +2y ^{2} -y -2\]

OpenStudy (fibonaccichick666):

In terms of h,k,r

OpenStudy (anonymous):

standard form is\[(x-h)^{2} + (y-k)^{2} = r ^{2}\]

OpenStudy (fibonaccichick666):

alright good, so first step, tell me how you found the center

OpenStudy (fibonaccichick666):

what did you do?

OpenStudy (anonymous):

h=-2 and k=1/4 is what I got for completing the squares and getting it into standard form and that is right. I was told my radius was wrong. I have tried 13 times and gotten different answers each time. Im on a program that's practice problems for calculus and they wont tell me why its wrong.

OpenStudy (fibonaccichick666):

ok, so walk me through what you did so I can help you figure out the mistake

OpenStudy (anonymous):

give me a minute to type it up on here my notes are a disaster

OpenStudy (fibonaccichick666):

alright, you can do multiple things if you want

OpenStudy (anonymous):

okay, in all honesty, I started this problem earlier and stop for a while and moved on because I was getting frustrated. now I cant find the notes, but what I got was correct, I just cant remember how I got those answers, im drawing a complete blank.

OpenStudy (kohai):

Graph it, and then the radius is the point smack in the middle

OpenStudy (anonymous):

wouldn't that be the center point not the radius?

OpenStudy (anonymous):

and I can't graph it without the radius

OpenStudy (kohai):

Yes you can, I just did. When you find the center, you'll find the radius

OpenStudy (fibonaccichick666):

ok, so one, calculators are not allowed in most colleges, I strongly recommend learning the written approach 2 just verbalize and tell me for now the process and we can do it together

OpenStudy (kohai):

I don't believe that for a second. It's good to know the written way, but I simply don't believe that calculators aren't allowed. This might help you visualize.

OpenStudy (fibonaccichick666):

I was not allowed a calculator past 12th grade and I am now considering a phd in math. It's true. We are not allowed calculators

OpenStudy (anonymous):

we have to take all pieces with variables on the left side, get numbers on the right. Then complete the two squares on the lefthand side for x variables and then y variables. then factor and you have the two squares that make h=-2, and k=1/4

OpenStudy (fibonaccichick666):

alright, now that is one way, good, but my question is what does that give you on the right where you have your numbers?

OpenStudy (anonymous):

I had -1, but I don't think that is right and im not quite sure how I got that

OpenStudy (fibonaccichick666):

ok, so you are probably right

OpenStudy (fibonaccichick666):

it probably shouldn't be negative one, but let's generalize one step further. What should the -1 represent?

OpenStudy (fibonaccichick666):

(in theory)

OpenStudy (anonymous):

radius?

OpenStudy (fibonaccichick666):

there is the key, it represents the radius squared

OpenStudy (fibonaccichick666):

so the reason -1 is weird is because you then have an imaginary radius (\(\sqrt{-1}=i\))

OpenStudy (anonymous):

but that cant be right. I tried showing it was the squareroot of -1 and it said that's impossible, then I tried putting in "I" but it wont allow me

OpenStudy (fibonaccichick666):

right, probably because when you completed the square, you forgot to move the extra numbers to the right side

OpenStudy (anonymous):

I know that the number on the right side represents radius squared and to take the square root of it but I keep getting wrong answers, ive tried multiple times, can you please show me what you get for radius squared

OpenStudy (fibonaccichick666):

so let's think, half the work is done for us since we know the h,k values

OpenStudy (fibonaccichick666):

we need to first look at the completing the square bit

OpenStudy (fibonaccichick666):

\[2x^2+8x+2y^2−y−2=0\] This is our given correct?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

alright so my first step is to get rid of the two's on the x^2 and y^2. Can you think of why?

OpenStudy (anonymous):

because it wouldn't be possible to complete the square. It would make it this,\[2(x ^{2} +8x) 2(x ^{2}-y) = 2(-1)\]

OpenStudy (fibonaccichick666):

first part correct equation not quite but you got the idea

OpenStudy (fibonaccichick666):

so, in order to rid myself of the 2's. I am going to multiply both sides by _____ to get \(x^2+4x+y2−\frac{1}{2}y−1=0\)

OpenStudy (anonymous):

-2

OpenStudy (fibonaccichick666):

no, we are going to multiply by \(\frac{1}{2}\) can you tell me why?

OpenStudy (fibonaccichick666):

and remember we do the same thing to both sides

OpenStudy (anonymous):

to get rid of the two, its pretty much the same as dividing it by two

OpenStudy (fibonaccichick666):

yea, can you show me how to do that though, in an eq so I know you get the concept

OpenStudy (anonymous):

what do you mean show you how to do that, you already did it up there?

OpenStudy (fibonaccichick666):

Can you do it step by step? I skipped two steps

OpenStudy (anonymous):

which steps did you skip because I don't see them. to be able to complete the square we had to get rid of the 2's therefore you multiplied both sides to get the equation you got above. there was nothing skipped?

OpenStudy (fibonaccichick666):

yea, I skipped the actual multiplication steps. Those are what I need to see before we continue

OpenStudy (fibonaccichick666):

so first step write in the mult. second simplify third simplify more (now that I think there are actually 3 steps)

OpenStudy (anonymous):

\[2x ^{2}(1/2) +8x(1/2) +2y ^{2}(1/2) -y(1/2) -2(1/2) = x ^{2}+4x +y ^{2} - 1/2y -1\]

OpenStudy (fibonaccichick666):

ok so I'm gonna call you on a technicality that will get you into trouble later, what does that all equal?

OpenStudy (anonymous):

\[x ^{2}+4x +y ^{2}-1/2y = 1\]

OpenStudy (fibonaccichick666):

yes and no, not what I was fishing for but correct. So you cannot forget to write : \[2x2(1/2)+8x(1/2)+2y2(1/2)−y(1/2)−2(1/2)=\color{red}{0\circ\frac{1}{2}}\] otherwise it is incorrect

OpenStudy (fibonaccichick666):

it seems silly but trust me, teachers bust your you know what for those small inconsistencies

OpenStudy (anonymous):

okay, understandable

OpenStudy (fibonaccichick666):

alright so now that I know you can do the multiplication let's do the complete the evil square that we all detest. So key thing, keep your eq =0 to start. it makes life easier in the end

OpenStudy (fibonaccichick666):

We now have: \[x^2+4x+y^2−1/2y-1=0\]

OpenStudy (fibonaccichick666):

so start with the y's we know we need a -1/4 right?

OpenStudy (anonymous):

\[x ^{2} +4x +2 \] and \[y ^{2} -1/2y +1/8\]

OpenStudy (fibonaccichick666):

uhm yea, that would be what we need, but now look at non-variable number that we have vs the ones we need

OpenStudy (fibonaccichick666):

we have -1 but we need a 1/8 and a *correction* 4

OpenStudy (fibonaccichick666):

so we need to add zero in order to have those numbers, how can we add zero?

OpenStudy (anonymous):

I don't quite understand what you're saying,

OpenStudy (fibonaccichick666):

so this is a trick that we need to use. You sort of just have to memorize it's applications so you can recognize when you need it. what is \[\frac{1}{8}+\frac{-1}{8}=?\]

OpenStudy (anonymous):

0

OpenStudy (fibonaccichick666):

yup so if we add that and the respective for 4 to both sides, we don't mess with our nswer because 0+anything =anything right?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

alright, so what we shall do is just what I said above, let's show it in math speak, then you can finish the problem for me, ok?

OpenStudy (anonymous):

okay

OpenStudy (fibonaccichick666):

\[(x^2+4x+4)-4+(y^2−1/2y+\frac{1}{8})-\frac{1}{8}-1=0\]

OpenStudy (fibonaccichick666):

is the math speak for what I just said. To make it easier, I threw in some parenntheses since addition is commutative

OpenStudy (fibonaccichick666):

and associative

OpenStudy (anonymous):

\[(x+2)^{2} + (y-1/4)^{2} = 1+4+1/8 = 5and1/8\]

OpenStudy (fibonaccichick666):

yup, now what does r=?

OpenStudy (anonymous):

\[\sqrt{41/8}\]

OpenStudy (fibonaccichick666):

mhmm

OpenStudy (fibonaccichick666):

do you understand?

OpenStudy (anonymous):

Yes, i understand but my program says this is wrong

OpenStudy (anonymous):

@Destinymasha this is the question if you got my message and have the time

OpenStudy (anonymous):

can someone please check the answer we have gotten above. the program i am using says that our radius is incorrect. if you could tell me why and how o get to the correct answer, that would be amazing! i really just want to understand why i am getting this wrong everytime i try.

OpenStudy (anonymous):

@Hero could you possibly take a look

OpenStudy (anonymous):

@SolomonZelman could you maybe take a look at this?

OpenStudy (anonymous):

@beccaboo333 could you take a look at this please?